# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

### v2

@@ 5.2.1. Apparatus
The baseline condition presented a side-by-side interface with a text editor and GPT-4o conversational panel, enabling both manual editing and LLM-assisted text/image generation. Participants completed two story-writing tasks (Appendix A.1), each extending a given story beginning into a 300–500 word draft. Tasks were counterbalanced across conditions (Baseline vs. Vistoria).
@@ 5.2.2. Study Procedure
In each condition, participants first focused on worldbuilding and idea exploration (20 min) and then on refining and improving the story (20 min). We divided the writing task into two phases (exploration and refinement) to prevent participants from prematurely committing to a single storyline and to reduce fixation, thereby encouraging broader ideation before focused improvement (Suh et al., 2024).
After each condition, participants completed surveys including NASA-TLX (Hart and Staveland, 1988), Creativity Support Index (CSI)(Cherry and Latulipe, 2014), self-perceived ML experience(Choi et al., 2024; Wu et al., 2022), and (for Vistoria only) perceived feature usefulness. All instruments used 7-point Likert scales. Following both conditions, participants completed a 15–20 minute semi-structured interview.
All sessions were video-recorded via Zoom. We collected system logs, final story drafts, canvas artifacts, image–text pairs, and interview transcripts for analysis.
@@ 5.2.3. Data analysis
First, interview transcripts were independently open-coded by two authors following an inductive, grounded theory–informed process (Thornberg, 2012); initial codes were iteratively refined into a shared codebook, with themes developed through constant comparison. Disagreements were resolved via negotiated agreement, and analytic memos were maintained for transparency.
Second, interaction data were analyzed through structured video coding by two authors and aligned with system logs on an event-by-event basis. We examined the sequences of function use and compared exploration patterns across the baseline and our condition. To characterize participants’ divergent–convergent behaviors during the creative process, for each task, we reconstructed exploration structures by defining Directions as top-level trajectories toward a goal, Branches as the diversity of possibilities generated within a direction, and Depth as the mean number of iterative steps within each branch to compare the exploration across the baseline and our condition.
Finally, for survey measures, we conducted paired-sample t-tests under the assumptions of normality and homogeneity of variance. When assumptions were violated, we used the Wilcoxon signed-rank test.
@@ 6. Study Results
Transparent of ML model Transparency 4.250 1.865 3.417 2.234 0.2015
Controllability 4.750 1.357 4.833 1.586 0.8664
Sense of Collaboration 5.583 1.505 4.333 1.670 0.0206 *
Support for Thought Process 5.333 1.371 4.750 1.960 0.3172
Match to goal 5.500 0.905 4.833 1.337 0.1201
@@
6.1. Strategies for Integrating Multimodality in Creative Writing
To address RQ1, we identified four key interaction strategies employed when using Vistoria.
First, participants leveraged multimodal elements to externalize and interconnect different branches of their ideas. Second, they utilized image-text pairs as catalysts to generate richer and more dynamic narrative descriptions. Third, participants employed visual instruments as a mechanism to drive pivotal shifts in their storytelling. Finally, they treated images as both checkpoints and traces, enabling them to maintain narrative coherence and facilitate recall throughout the story development process.
6.1.1. Use Multimodality to Support Expressing and Connecting Ideas
Log data (Figure 6) and observations (Figure 13) reveal that participants followed a consistent creative workflow when using the system. The process typically began with multimodal content generation, where participants combined generation functions across multiple modalities to visualize their intended content. During this exploratory phase, participants iteratively collected and refined story elements through generated cards(images-text pairs), at the same time, frequently modifying text in the editor or regenerating images based on updated text.
As participants explored multiple visualization directions, they employed the Collage function to combine existing canvas content and discover new narrative possibilities, or used visual instruments (Lasso, Filter, Perspective Shifts) to simultaneously modify text and images. This process involved continuous comparison with LLM outputs to determine desired versus undesired elements (P1, P2, P4, P8, P10). Participants systematically highlighted useful images and text for final translation into the text editor, often using Cluster functions to summarize character or object settings for incorporation into their final writing.
Multimodal Expression Enhances Creative Intention. The multimodal integration enhanced participants’ ability to convey creative intentions. As shown in Figure 6 (b), text served as the primary medium but was consistently supplemented with sketches and images for context, yielding significantly higher expressiveness ratings than the baseline ($M_{\text{Vistoria}}=6.08$ vs. $M_{\text{Baseline}}=4.33$, p = .023) (Table 2). P6 captured this benefit: “Even though I’m not very good at drawing, when I type some texts and sketch a few rough images, the cards with story and images generated can really capture the scene I have in mind.” (Figure 11). Nearly all participants (P2, P6, P10, P11) stressed that combining sketches with text and images aligned outputs more closely with their creative intentions. P7 illustrated this with a concrete case: she sketched a rough flower and archway, then added text specifying a glowing flower and a Gothic gate. The system fused the spatial layout from the sketch with textual details to generate multiple fitting images. She highlighted the unique value of sketching: “Drawing is pretty cool. In GPT, I want to draw a mental image, but GPT cannot… the geometry is always different from what’s in my head.“ (Figure 7). This suggests that multimodal expression enabled participants to externalize their mental imagery and refine it into concrete, shareable representations, bridging the gap between vague internal visions and precise creative outputs.
Figure 7. Multimodal Expression. Example where sketch structure + textual details jointly yield multiple relevant images; participants valued sketches to visualize spatial layout beyond what text-only tools could provide.
@@ 6.1.1. Use Multimodality to Support Expressing and Connecting Ideas
Collage Function Enables Creative Recombination. The Collage function emerged as a central strategy for creative recombination, receiving positive usability ratings (M=5.33, SD=1.5) (Figure 8). All participants used collage to merge extracted objects or scenes from different images, building connections across disparate elements. For instance, when P6 generated a scene of Maya entering a castle, he envisioned a larger structure with taller stairs. He sketched a bigger castle and mountain while expressing his intention through text, resulting in a generated card that matched his vision and was directly incorporated into his story (Figure 9). P11 articulated the creative freedom this technique provided: “This technique doesn’t limit me; I can create abstract or non-abstract sketches, and I can incorporate whatever I want.” This multimodal recomposition enabled participants to quickly express envisioned scenes (P2) and provided “more freedom to envision and create the story” (P10). These practices highlight that collage is not merely a usability feature but a catalyst for multimodal recomposition, enabling participants to externalize, reconfigure, and expand their mental imagery into coherent narrative possibilities that text or images alone could not achieve.
Furthermore, multimodal input also increased participants’ immersion with Vistoria compared to the baseline ($M_{\text{Vistoria}}=4.92$ vs. $M_{\text{Baseline}}=2.75$, p = .0006) (Table 2). This heightened immersion stemmed from the playful, enjoyable nature of multimodal creation and the emotional reinforcement of freely expressing intentions across modalities. As P1 reflected: “Using the tool feels more like… doing something on a whiteboard or a big piece of paper, where you can do almost anything—it’s very free and interesting.” Together, multimodal inputs and outputs fostered a more engaging and emotionally resonant creative experience.
Figure 8. The perceived usefulness of each key functionality in Vistoria
6.1.2. Instrumental Interaction Tools Drive Narrative Development
Analysis of interaction logs (Figure 6 and 13) shows that participants often engaged with visual instruments in the mid-phase of creation, and these instruments directly shaped how stories were developed and revised.
@@ 6.1.2. Instrumental Interaction Tools Drive Narrative Development
The Perspective Shift instrument, which altered both the image and the story perspective, was rated among the most useful features (M=5.92, SD=1.56) (Figure 8). P8 described how shifting perspectives changed the story direction: “Adopting the water’s viewpoint anthropomorphized the water spirit and introduced a regretful undertone that established the story’s emotional framework. Changing the perspective of the voice of the story added personal sentiment and changed the development direction, inspiring new writing possibilities.” P5 also experimented with this feature, incorporating a first-person voice (“I didn’t expect this to be so heavy!”) adopted from the system-generated segments into her third-person story (Figure 10 (a)). Perspective Shift allowed authors to flexibly reconfigure narrative viewpoint and voice, surfacing new emotional framings and redirecting story trajectories without disrupting their ongoing writing flow.
Lasso as a granularity controller for local-to-global rewriting.
The Lasso instrument was also valued as useful (M=5.0, SD=1.28) (Figure 8) for enabling writers to zoom between different narrative scales. P8 emphasized, “You can write in different scales, especially when you use the Lasso tool, in which you can extract out that specific detail, so [the story] generated in the card is more heterogeneous on the specific point.”(Figure 10 (b)) This reflects how the lasso tool enables a narrative “zoom” functionality, allowing writers to oscillate between macro-level story development and micro-level detail refinement within a single interface. Similarly, P9 and P11 described how the lasso allowed them to deepen or refine specific text and emphasize the key points with more focused attention. In this way, the Lasso operates as a narrative instrument, turning macro-level edits into micro-level adjustments while preserving precision, enabling rollback, and sustaining fluent exploration.
@@
Filters as affective parameterization for tone alignment.
The Filter instrument shaped narrative emotion and tone by visually parameterizing affect. P11 noted, “The image provides the style, which influences my story’s tone and direction… After using the filter, I can’t determine tone from text alone—I need to choose between suspenseful or romantic, and the visuals help me decide.” Similarly, P9 observed, “Filter visualizes the emotions in my mind when creating stories… I created a sad mood with filters, and the girl suddenly turned into a sad face. It’s very easy to see what it’s doing.” (Figure 10 (c)). The immediate visual feedback fostered a more intuitive workflow, aligning emotional intent with visual representation. The immediate visual feedback aligned emotional intent with representation, streamlining tone-setting decisions.
Taken together, these tools acted as narrative instruments that transformed localized operations into meaningful shifts in viewpoint, scale, and tone. They enhanced authors’ control, amplified expressiveness, and sustained a fluid perception–action loop throughout the creative process.
Figure 10. (a) Perspective Shift was rated highly useful (M=5.92); changing the image viewpoint also reframed narrative voice and redirected story development, as participants reported; (b) Filters (M=4.25) synchronize mood and style across media—applying a visual filter also rewrites the associated text to match the intended emotional tone, ensuring stylistic coherence; (c) Lasso tool (M=5.0) enabled writers to focus at different narrative scales by extracting or isolating elements to steer local and global edits.
6.1.3. Structural Mapping between Images-Text Paris as a Catalyst for Vivid Story Writing
Participants emphasized that when text and images appeared together, images served as direct references during writing, easing detailed description and enriching narrative depictions. This integration worked through two cognitive mechanisms that supported vivid story construction.
Images-Text Paris as cognitive scaffolding for unfamiliar scenarios. Visual references helped participants imagine actions or settings beyond their lived experience. As P11 reflected, “I’m more comfortable describing conversations and overall scenarios than people’s actions—especially kids, who might have more interesting reactions. Visualizing with text descriptions helps me imagine unfamiliar actions, which I can then use in my story.” Here, the image acted as a structural anchor by aligning the visual cues of children’s postures and expressions with possible textual descriptions. In other words, structural mapping between what was seen (visual structure) and what could be narrated (linguistic structure) scaffolded richer and more plausible action descriptions.
Images-Text Paris enabling descriptive language translation. Participants also converted visual elements into vivid prose, effectively mapping structures across modalities. As P5 noted, “Looking at an image makes it easier to kind of, like, describe what something is supposed to look like or feel like… so when you describe it to [readers], and they can picture it for themselves in their head, that is what makes the story more interesting.” From observation, P5 closely described the final images in her writing and directly adopted text from the generated cards, integrating them into her final story. (Figure 12) This visual-to-textual translation was also evident in concrete writing outcomes. For example, P1 used visual cues from an image showing Claire touching a letter and adopted the descriptions in the text such as “Claire steadies the box on her lap” and “She runs her fingers over the letters, heartbeats echoing in her ears.” to form his final story. (Figure 11) Here, the structural alignment between gesture in the image and embodied description in text allowed P1 to carry emotional undertones across modalities, turning static imagery into dynamic, sensory-rich narrative moments.
Overall, structural mapping between visual and textual representations scaffolded descriptive richness, enhanced vividness, and supported schema formation for scenarios participants might otherwise have struggled to imagine.
Figure 11. Visual-to-text translation: writers turn visual cues into vivid prose that readers can picture; sample lines show how P1 used sensory-rich descriptions derived from an image-text pair.
6.1.4. Structural Mapping across Multiple Dimensions to Trace the Story
Participants used not only text but also images as visual checkpoints to trace story direction and development. Because images were more immediately perceptible than text, they helped participants align outputs with intended style or mood and, when viewed in sequence, detect progressions and gaps (P6, P7, P2). Within Vistoria, this process evolved into structural mapping: participants systematically compared visual–textual structures across dimensions, gaining multimodal visibility into narrative development.
This layered view created a relational space where main storylines and branching alternatives could be mapped against one another, allowing participants to perceive both continuity and divergence, thereby fostering coherence while supporting flexible redirection. As P12 noted, visuals alongside text set the mood and made it easier to recall earlier ideas, illustrating how structural mapping provided mnemonic scaffolding that supported continuity in storytelling. Compared to GPT’s linear text-only workflow, participants found that Vistoria’s visual traces better supported narrative management and reduced idea drift or loss (P12, P11, P2). P2 reflected on this benefit explicitly:
“Using the system, we started with a baseline picture and then came up with another picture… it was easy to trace the development of the story. If you get too far down that chain and don’t like it, you can just delete that node and go in a different direction. I liked visually being able to see the progress from generation to generation.”
Here, structural mapping across successive images enabled participants to align branches and selectively prune or redirect them. Such multi-dimensional mapping clarified narrative progression, improved recall and coherence, and supported flexible branching and gap detection, sustaining sensemaking throughout the evolving story.

### v3

@@ 5.2.1. Apparatus
The baseline condition presented a side-by-side interface with a text editor and GPT-4o (38) conversational panel, enabling both manual editing and LLM-assisted text/image generation. Participants completed two story-writing tasks (Appendix A.3), each extending a given story beginning into a 300–500 word draft. Tasks were counterbalanced across conditions (Baseline vs. Vistoria).
@@ 5.2.2. Study Procedure
In each condition, participants first focused on world-building and idea exploration (20 min) and then on refining and improving the story (20 min). We divided the writing task into two phases (exploration and refinement) to prevent participants from prematurely committing to a single storyline and to reduce fixation, thereby encouraging broader ideation before focused improvement (Suh et al., 2024).
After each condition, participants completed surveys including NASA Task Load Index (TLX) (Hart and Staveland, 1988) and Creativity Support Index (CSI) (Cherry and Latulipe, 2014). These surveys were chosen to be consistent with the standard measures employed in previous HCI system work on multimodal LLM-assisted ideation and storytelling (Choi et al., 2024; Suh et al., 2025; Darejeh et al., 2024). All surveys used 7-point Likert scales. After both conditions finished, participants also completed a 15–20 minute semi-structured interview. All sessions were video-recorded via Zoom. We collected system logs, final story drafts, canvas artifacts, image–text pairs, and interview transcripts for analysis.
@@ 5.2.3. Data analysis
For the qualitative interview data, we conducted an inductive, grounded theory-informed analysis (Thornberg, 2012). First, two authors independently performed open-coding on 33% of the data, generating an initial set of 30 distinct codes. The coders then met to compare code applications, resolve discrepancies through negotiated agreement, and refine the wording and boundaries of each code. Through several rounds of discussion, they reached full consensus on all coded segments and consolidated the initial codes into a shared codebook. Using the refined codebook, the two authors independently coded half of the remaining transcripts, meeting regularly to prevent coding drift and to determine whether newly emerging codes should be incorporated. Ongoing constant comparison within and across interviews was used to further refine relationships between codes. Finally, we clustered the codes into four higher-level themes that map onto our design goals. The final codebook is shown in Appendix A.6.
Second, interaction data were analyzed through structured video coding by two authors to quantify tool usage frequency and modality switching events, and researchers aligned them with system logs on an event-by-event basis. Finally, for survey measures, we conducted paired-sample t-tests under the assumptions of normality and homogeneity of variance. When assumptions were violated, we used the Wilcoxon signed-rank test. Given the sample size limitations, we treat these quantitative results as descriptive signals intended to triangulate with and support the qualitative themes.
@@ 6. Study Results
6.1. The Usability of Vistoria
To address RQ1 and RQ2, we analyzed how participants engaged with the designed instrumental operations (Lasso, Collage, perspective shift, and Filter) and the usage patterns. In addition, we incorporated quantitative survey results with qualitative data to assess how Vistoria affected participants’ workload.
6.1.1. The Usefulness of Instrumental Operations
Lasso as a granularity controller for local-to-global rewriting.
The Lasso instrument helped participants shift between narrative scales, supporting both macro-level story planning and micro-level detail refinement. As P8 noted, “You can write in different scales, especially when you use the Lasso tool, in which you can extract out that specific detail, so [the story] generated in the card is more heterogeneous on the specific point.” (Figure 7(c)). Similarly, P9 and P11 described using the Lasso to foreground key points and sustain more focused attention during revision. Together, these accounts suggest that the Lasso affords a narrative “zoom,” enabling users to isolate a local fragment for targeted rewriting while maintaining continuity with the broader narrative trajectory within the same interface.
Figure 7. (a) Perspective Shift changes image viewpoint to reframe narrative voice and redirect story development; (b) Filter synchronizes mood and style across media—applying a visual Filter also rewrites the associated text to match the intended emotional tone; (c) Lasso enables participants to operate at different narrative scales by extracting or isolating elements to steer local and global edits.
The figure provides examples of instrumental operations in use. It shows how Perspective Shift reframes narrative voice, how Filters synchronize emotional tone across images and text, and how the Lasso enables writers to operate at different narrative scales by isolating or extracting specific elements for local or global edits.
Collage Function Enables Creative Recombination.
@@ 6.1.1. The Usefulness of Instrumental Operations
Participants used Collage to merge extracted objects or scenes from different images, building connections across disparate elements. For instance, when P6 generated a scene of Maya entering a castle, he envisioned a larger structure with taller stairs. He sketched a bigger castle and mountain while expressing his intention through text, resulting in a generated card that matched his vision and was directly incorporated into his story. P11 articulated the creative freedom Collage technique provided: “This technique doesn’t limit me; I can create abstract or non-abstract sketches, and I can incorporate whatever I want.” (Figure 8). Collage thus serves not merely as a usability feature but as a catalyst for externalizing, reconfiguring, and expanding mental imagery into narrative possibilities unachievable through text or images alone.
@@
The Perspective Shift alters both the image and the story perspective to provide new directions for story development. P8 described how shifting perspectives changed the story direction: “Adopting the water’s viewpoint anthropomorphized the water spirit and introduced a regretful undertone that established the story’s emotional framework… changed the development direction, inspiring new writing possibilities.” P5 also experimented with this feature, incorporating a first-person voice (“I didn’t expect this to be so heavy!”) adopted from the system-generated segments into her third-person story (Figure 7 (a)). Perspective Shift allowed participants to flexibly reconfigure narrative viewpoint and voice, surfacing new emotional framings and redirecting story trajectories without disrupting their ongoing writing flow.
Filter as affective parameterization for tone alignment.
The Filter instrument shaped narrative emotion and tone by visually parameterizing affect. P11 noted, “The image’s style influences my story’s tone and direction. Text alone doesn’t reveal tone. The Filter’s visual changes help me decide the desired feeling when I need to choose between suspenseful or romantic approaches.” Similarly, P9 observed, “Filter visualizes the emotions in my mind when creating stories… I created a sad mood with Filters, and the girl suddenly turned into a sad face. It’s very easy to see what it’s doing and easy for me to describe later.” (Figure 7 (b)). The immediate visual feedback aligned emotional intent with text, streamlining tone-setting decisions to evaluate the usability.
Taken together, these instrumental operations transformed localized operations into meaningful viewpoints, scales, and tones. They appear to support participants’ intended operational precision and expressiveness, while potentially reinforcing the perception–action loop during the planning and translating of story writing.
@@
The figure demonstrates the Collage operation during story writing. Participants combine extracted objects and scenes from multiple cards, often mixing sketches with images, to specify visualization and explore new narrative ideas. The example highlights creative recombination as a mechanism for advancing story development.
6.1.2. Tradeoffs of Multimodal Image-Text Co-editing
Increased Workload.
While image–text co-editing may support aspects of narrative coherence and expressiveness, our results suggest that it can also introduce additional workload. NASA–TLX scores showed significantly higher mental demand (NASA-TLX: $M_{\text{Vistoria{}}}=5.16$ vs. $M_{\text{baseline}}=3.17$, p=.0000; Table 1) and physical demand (NASA-TLX: $M_{\text{Vistoria{}}}=4.67$ vs. $M_{\text{baseline}}=2.08$, p=.0002; Table 1), plus moderately higher frustration (NASA-TLX: $M_{\text{Vistoria{}}}=2.75$ vs. $M_{\text{baseline}}=1.75$, p=.0204; Table 1). Qualitative analysis reveals that this increased workload reflected the higher cognitive and physical effort of coordinating across modalities and actively curating outputs compared with the GPT baseline . Part of the mental load may have stemmed from first-time use—learning new image/text operations and switching between modalities (P1, P2, P5, P9, P10). Additional difficulty also arose from unfamiliarity with the canvas interface compared with the traditional GPT interface (P3, P10). Below we present the tension of Image-Text co-editing based on qualitative feedback.
The Cognitive Effort of Enhanced Sense of Agency. Nine participants (all except P2, P7, P12) felt they have higher sense of agency, and valued maintaining story control and felt they were directing the narrative through their sketches, inputs, and use of system tools (P4, P6, P10). However, P1, P3, P9, and P10 noted that they had to actively develop details within their own text, especially at the early stages, which could feel “a little bit frustrating” (P9). Unlike GPT, which could quickly produce long passages or propose questions to guide brainstorming (P10), the system requires participants to supply and elaborate on their own ideas before meaningful generation occurs, potentially leading to a higher mental workload. This shift demanded more cognitive and physical effort: even though sketching and annotation helped externalize mental imagery, participants noted that it felt more demanding than simply inputting and receiving GPT’s ready-made text (P1, P2, P11). Thus, a stronger sense of agency may come at the cost of a higher workload.
Validating Ideas Rather than Generating Them. Some participants noted that the system’s strengths lay in validating or expanding existing concrete mental images rather than generating new and abstract directions (P1, P11). As P1 explained, “When I have a vague impression in my mind, I tend to generate some image-text pairs. But sometimes, once the visual appears, it fixes my imagination in a certain way in my mind, and I can no longer imagine other possibilities. In contrast, only plain text can inspire limitless imagination.” This reveals a tension: images act as concrete anchors that aid detailed development, yet their representational specificity can induce fixation by prematurely crystallizing fuzzy concepts and narrowing exploration. Similarly, P10 noted that the baseline GPT condition was superior at breaking down initial story points and directly providing additional suggestions and directions, whereas the Vistoria system primarily served to elaborate or diverge from existing visions the user already has. This suggests this multimodal approach may be valuable for participants with partially formed concepts, though potentially less helpful during the open-ended phases of ideation when abstract exploration is more important than visual specificity.
Externalization Frees Cognitive Space. Although participants reported experiencing higher mental workload, several accounts suggest that multimodality may have supported a more efficient allocation of attention (P4, P7). As P7 described, “By highlighting and collaging, I externalized formed ideas into image–text pairs, clearing mental space to pre-plan the next line and concentrate on the next plot beat.” It indicates that the creation process has the potential to externalize fleeting ideas, preserve sensory and spatial detail, and reduce information loss when translating imagination into concrete artifacts.
6.2. Creativity Support of Vistoria
To address RQ3 and RQ4, we examine how Vistoria supports intention expression through multimodal input, facilitates the ideation process through divergent exploration, as well as how image–text alignment design contributes to narrative development. We also describe how this workflow potentially enhances participants’ sense of agency and ownership.
Figure 9. Multimodal Expression. Example where sketch structure and textual details jointly yield multiple relevant images; participants valued sketches for visualizing spatial layout beyond what text-only tools could provide.
The figure shows an example of multimodal expression where rough sketches and textual descriptions jointly guide image generation. The resulting images reflect the spatial layout and visual structure conveyed by sketches that would be difficult to express using text alone, supporting more precise externalization of mental imagery.
6.2.1. Enhancing Intent Expression Through Multimodal Input.
As shown in Appendix Figure 13 (b), text served as the primary medium for generation but was consistently supplemented with sketches and images to provide more spatial information. Eight participants (P1, P2, P3, P6, P7, P9, P10, P11) stressed that combining sketches with text and images aligned outputs more closely with their creative intentions, yielding significantly higher expressiveness ratings than the baseline ($M_{\text{Vistoria}}=6.08$ vs. $M_{\text{Baseline}}=4.33$, p = .023; Table 1).
P6 captured this benefit: “Even though I’m not very good at drawing, when I type some texts and sketch a few rough images, the cards with story and images generated can really capture the scene I have in mind.”. P7 illustrated this with a concrete case: she sketched a rough flower and archway, then added text specifying a glowing flower and a Gothic gate. The system fused the spatial layout from the sketch with textual details to generate multiple fitting images. She highlighted the unique value of sketching: “Drawing is pretty cool. In GPT, I want to draw a mental image, but GPT cannot… the geometry of GPT-generated image is always different from what’s in my head.” (Figure 9). This suggests that multimodal expression enabled participants to externalize their mental imagery and refine it into concrete, shareable representations, bridging the gap between vague internal visions and precise outputs.
6.2.2. Bottom-up Creation to Support Divergent Exploration
As shown in Appendix Figure 12 and Appendix Table 2, participants demonstrated greater breadth and depth of exploration using Vistoria and produced visibly more divergent narrative structures compared to the linear outputs of the GPT baseline. The Vistoria canvas functioned as an exploratory space where participants pursued multiple storylines in parallel without linear constraints.
Rather than committing immediately to single narratives, participants typically generated multiple alternatives in early phases. As P8 observed: “GPT workflow is more streamlined… top-down. Using the system feels more bottom-up. You are open to possibilities, and then you choose one way to go deep, so there’s not a finite result and more possibilities being explored.”
The exploratory nature of Vistoria was also described as playful and enjoyable (P1, P7). As P1 reflected: “Using the tool feels more like doing Collage or drawing on a whiteboard or a large sheet of paper, where you can do almost anything—it’s very free and interesting.” This experiential quality aligns with the higher immersion reported for Vistoria compared to the baseline (CSI: $M_{\text{Vistoria}}=4.92$ vs. $M_{\text{Baseline}}=2.75$, p = .0006; Table 1).